# Slow-phonon control of spin Edelstein effect in Rashba $d$-wave altermagnets

### v1

@@ IV Results and discussion
A key point worth asking is whether phonon-induced suppression or complete vanishing of the spin Edelstein effect can be effectively harnessed to control spin transport in practical spintronic applications. Our analysis indicates that controlled phonon scattering offers precisely this opportunity. When phonon interactions are moderate, they act to dampen excess spin polarization, thereby minimizing spin-related dissipation in spin-orbit torque devices. In contrast, under strong phonon-mediated spin depolarization is central to the operation of spin logic circuits [65]. Spin-polarized states represent the binary logic levels required for data encoding and signal transmission, while spin depolarization plays an equally critical role by resetting logic units, erasing residual spin memory, and electrically isolating different parts of the circuit to prevent spin leakage.
Since the (de)polarization occurs primarily in the dominant transverse susceptibilities, and given that $\chi_{\ell j}\neq-\chi_{j\ell}$ but behaves similarly to $\chi_{j\ell}$, we henceforth focus on $\chi_{xy}$ and analyze its evolution across the parameter space in the presence of EPC.
@@ IV.1 Altermagnetic order effect
Next, we analyze how altermagnetism affects the persistence of spin Edelstein depolarization, allowing us to contrast the behavior with and without an intrinsic spin-split band structure. Figure 5(a) shows the dependence of the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ on the EPC strength $g$ for different values of the staggered lattice potential parameter, or equivalently, the altermagnetic order $\beta$. For $\beta=0$, the system also exhibits a gradual suppression of $\chi_{xy}/\chi_{0}$ as $g$ increases because it is well established that RSOC alone generates a finite spin Edelstein polarization as a direct consequence of inversion-symmetry breaking. Importantly, the depolarization transition, signaled by the zero of $\chi_{xy}/\chi_{0}$, still occurs even in the absence of altermagnetism.
@@
Although the model inherently predicts an isotropic spin Edelstein polarization at $\beta=0$, corresponding to antisymmetric spin susceptibilities, we explicitly validate this behavior by examining the electronic band structure and DOS as in the preceding analyses. Figure 5(b) offers further insight by displaying the electronic DOS for $\beta=0$ and $g=0.2$ eV/Å, along with the corresponding electronic band structure $\mathcal{E}_{\vec{k}}$ along the high-symmetry directions $Y\leftarrow\Gamma\to X$ (inset). As illustrated, the DOS along both the $X$ and $Y$ directions exhibits identical features at $\beta=0$, confirming that in the absence of altermagnetic order, the induced spin Edelstein (de)polarization remains highly isotropic. By contrast, the introduction of a finite altermagnetic order parameter $\beta$, see Fig. 2, breaks this isotropy by inducing anisotropic spin-splitting in the band structure, thereby rendering the spin Edelstein response directionally dependent. This anisotropy is a crucial attribute for spintronic applications, where control over the spin polarization along specific crystallographic directions is often desired.
As $\beta$ increases, the induced spin polarization decreases and the depolarization point shifts slightly toward larger critical EPC strengths $g_{\rm c}$, indicating that the altermagnetic order delays the onset of depolarization. This was already confirmed by Fig. 3(a). Physically, this can be attributed to the enhanced band structure induced by the combined effects of EPC and altermagnetic order, which strongly suppresses the intraband and interband contributions to the susceptibility.
It should be noted that in conventional spin-orbit coupled systems, the Edelstein effect produces a nonequilibrium spin polarization that is subsequently reduced by relaxation processes, e.g. Dyakonov–Perel-type dephasing from momentum scattering in $\vec{k}$-dependent spin-orbit fields [82, 83] or Elliott–Yafet scattering from impurity and phonon interactions [84, 85]. These mechanisms dynamically limit the lifetime of spin accumulation but do not break the underlying asymmetry in induced polarizations. By contrast, the Edelstein (de)polarization in altermagnets breaks this asymmetry: opposite spin textures on different parts of their Fermi surface change upon Brillouin-zone integration, yielding anisotropic (de)polarizations even in the absence of scattering. This makes the effect intrinsic and symmetry-driven, rather than extrinsic and relaxation-driven.
@@ IV.2 Phonon-altermagnetic order coupling effect
As the next investigation, we examine how EPC to the altermagnetic order, $\tilde{g}/g$, affects the persistence of spin Edelstein depolarization. It is observed, in Fig. 6(a), that for $\tilde{g}/g<1$, the rate of polarization suppression depends only weakly on $\tilde{g}/g$. Larger values of the ratio (but still below 1) lead to a slower increase in $\chi_{xy}/\chi_{0}$ and slightly shift the zero to larger $g$, implying that a stronger coupling to the staggered potential slightly enhances the susceptibility renormalization. At $\tilde{g}/g=1$, there is a strong shift of the spin bands, such that a much larger EPC is required to suppress the spin Edelstein polarization, as the intraband and interband transitions still contribute significantly to the susceptibility. This, in turn, means that stronger $\tilde{g}/g$ stabilizes the transverse polarization more efficiently.
@@
We further highlight that in the absence of explicit EPC to the staggered potential, depolarization still occurs. However, in Fig. 6(b) at $g=0.2$ eV/Å, our DOS analysis and band structure (inset of Fig. 6(b)) reveal that the induced polarization and depolarization remain isotropic at $\tilde{g}/g=0$, since the DOS is identical along the $X$ and $Y$ directions of the Brillouin zone. The parameters $\tilde{g}$ and $\beta$ are intrinsically linked, such that changes in one necessarily influence the other, reflecting their mutual dependence within the system. Similar to the case of $\beta=0$, such isotropy is undesirable for spintronic applications, where directional control of spin responses is essential. In contrast, once the ratio $\tilde{g}/g$ is introduced, the polarization and depolarization become anisotropic (see Fig. 2), providing the anisotropic spin responses that are precisely what spintronics aims to exploit.
@@ IV.3 Rashba spin-orbit coupling effect
IV.4 Electron and hole doping effects
Finally, we examine the case of the doping effect, in which the chemical potential is shifted away from the Fermi energy ($\mathcal{E}=0$ eV). Figure 8 demonstrates the influence of the chemical potential $\mu$ on the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ as the EPC strength $g$ is varied.
@@ IV.4 Electron and hole doping effects
For positive chemical potentials ($\mu=+0.02$ and $0.04~\mathrm{eV}$), corresponding to electron doping, the behavior shown in Fig. 8(b) differs significantly. The susceptibility curve exhibits a nonmonotonic dependence on $g$, characterized by an intermediate enhancement before eventually decaying. The peak in the susceptibility can be understood in the same way as in Fig. 7(b): for $\mu>0$ and $g>0.15$ eV/Å, an increasing number of states participate in intraband and interband transitions, leading to a temporary increase in susceptibility. Beyond this point, as the number of available states diminishes, the susceptibility begins to decrease. Thus, a strong particle-hole asymmetry appears in the Edelstein response under finite doping.
@@ V Experimental perspective
Before concluding, we comment on the experimental feasibility of our predictions. A promising route to experimentally probe the proposed effects is to interface a 2D $d$-wave altermagnet—such as thin films of KV2Se2[86], RbV2Te2[87], RuO2 (noting that the altermagnetic character of RuO2 remains under active investigation) [88, 89, 14, 90], and $\kappa$-Cl [91]—with an appropriate substrate at low temperatures to induce lattice vibrations in the altermagnet layer. Crucially, since the underlying phonon-mediated mechanism requires only a linear coupling between lattice displacements and spin states, the proposed effect should be realizable across a broad class of altermagnetic thin films. Phonons at nominally zero temperature can also be generated with inelastic neutron scattering providing a direct route: incident neutrons transfer both momentum and energy to the lattice, populating phonon states even at zero temperature [92, 93].
Electrostatic control over the RSOC can be achieved using top and bottom gate electrodes separated by high-quality dielectric layers. Complementary to phonon generation and gating, time- and angle-resolved photoemission spectroscopy enables direct visualization of spin textures, band splittings, and symmetry-breaking phenomena [94]. Additionally, the Edelstein effect can be probed in Hall-bar geometries through electrical transport measurements.
@@ VI Summary and outlook
In this work, we have examined how slow lattice vibrations shape spin responses in $d$-wave altermagnets, systems that combine vanishing net magnetization with spin-split electronic bands. By formulating the problem within a Rashba continuum model coupled to slow phonons through a Holstein-type interaction, and treating the EPC at the static-Holstein level, we have tracked the evolution of the Edelstein effect over a broad range of parameters. The Kubo formalism provided access to distinguish the specific contributions of intraband and interband processes.
@@
A central outcome of our study is that increasing EPC gradually suppresses the Edelstein spin polarization. Once the coupling exceeds a critical threshold, the system undergoes a spin Edelstein depolarization. At this threshold, the spin polarization collapses to zero permanently, signaling a qualitative restructuring of the spin channels. This regime is characterized by the disappearance of both intraband and interband transitions at zero Fermi energy. The (an)isotropic character of the suppressed polarization and depolarization strongly depends on the presence of altermagnetism. Lacking altermagnetic order, the effect is isotropic, whereas the interplay with altermagnetism produces pronounced anisotropy and eliminates the usual antisymmetry between spin susceptibilities. This finding suggests that altermagnetism is a key ingredient in realizing directionally selective spin responses, thereby enhancing prospects for spintronic applications. Our parameter-space analysis further revealed that the onset of depolarization is tunable. The critical EPC shifts with the strength of coupling to a staggered lattice potential and the Rashba interaction. Doping also plays a decisive role: moving the chemical potential away from the zero Fermi energy modifies the threshold EPC for depolarization. Negative chemical potentials (hole doping) reduce the susceptibility and promote faster depolarization, whereas positive chemical potentials (electron doping) enhance the susceptibility and delay the onset of depolarization.
In practice, phonons can be controlled through approaches such as strain engineering or substrate selection. The selective excitation of specific vibrational modes modulates the EPC and local crystal fields, thereby influencing the induced spin polarization. Thus, our study shows that moderate phonon scattering can be utilized to regulate spin-related dissipation in spin-orbit torque devices by attenuating excess spin polarization. In contrast, when strong phonon-mediated depolarization dominates, it enables efficient clearing of residual spin information, reinitialization of logic elements for subsequent operations, and electrical isolation of different circuit sections to prevent spin leakage, thereby providing a practical route for on-demand switching between spin-polarized and spin-neutral states. Together, these mechanisms provide complementary strategies for enhancing the efficiency and design flexibility of emerging spintronic technologies.
@@
Looking ahead, several intriguing directions emerge for future investigations. First, exploring the interplay of phonon-induced spin depolarization with other types of spin-orbit interactions, such as Dresselhaus or proximity-induced effects, could reveal richer anisotropic spin responses. Second, extending these studies to multilayer altermagnets or heterostructures may enable enhanced tunability through interlayer coupling and interface engineering. Third, time-resolved experiments, including ultrafast optical or terahertz pump-probe techniques, could probe the dynamical evolution of the Edelstein effect under controlled phonon excitations, providing direct verification of our predictions. Fourth, combining phonon engineering with strain, gating, or substrate manipulation could allow precise control over spin polarization, opening pathways for optimized spintronic device design. Finally, extending these concepts to other material classes, such as topological insulators, 2D magnets, or Rashba systems with strong electron correlations, may reveal universal features of phonon-controlled spin responses. We leave these topics to be explored in future work.
@@ Acknowledgments
M. Y. would like to acknowledge useful discussions with Ulrich Zülicke at the early stages of this work. M. Y. and J. K. F. were supported by the Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Contract No. DE-FG02-08ER46542 for the formal developments, the numerical work, and the writing of the manuscript. J. K. F. was also supported by the McDevitt bequest at Georgetown University. J. L. was supported by the Research Council of Norway through its Centres of Excellence funding scheme Grant No. 262633 and Grant No. 353894.

### v2

@@ IV Results and discussion
In conventional Rashba systems, the Edelstein response is strictly antisymmetric, $\chi_{\ell j}=-\chi_{j\ell}$. In altermagnets, this antisymmetry is generically broken by momentum-dependent spin splitting and its phonon-induced renormalization. The resulting anisotropy can be tuned continuously via the EPC strength, altermagnetic parameter $\beta$, staggered phonon coupling $\tilde{g}$, Rashba interaction $\lambda_{\rm R}$, and carrier doping, providing a direct route to control direction-dependent spin responses. Importantly, varying these parameters affects the magnitude rather than producing qualitatively different behavior between the $xy$ and $yx$ components. Since depolarization occurs mainly in the dominant transverse susceptibilities, and $\chi_{\ell j}\neq-\chi_{j\ell}$ yet behaves similarly to $\chi_{j\ell}$, we focus on $\chi_{xy}$ and analyze its evolution across the parameter space in the presence of EPC.
@@ IV.1 Altermagnetic order effect
Next, we analyze how altermagnetism affects the persistence of spin Edelstein depolarization, allowing us to contrast the behavior with and without an intrinsic spin-split band structure. Figure 5(a) shows the dependence of the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ on the EPC strength $g$ for different values of the staggered lattice potential parameter, or equivalently, the altermagnetic order $\beta$. For $\beta=0$, the system also exhibits a gradual suppression of $\chi_{xy}/\chi_{0}$ as $g$ increases because it is well established that RSOC alone generates a finite spin Edelstein polarization as a direct consequence of inversion-symmetry breaking. Importantly, the depolarization, signaled by the zero of $\chi_{xy}/\chi_{0}$, still occurs even in the absence of altermagnetism.
@@
Although the model inherently predicts an isotropic spin Edelstein polarization at $\beta=0$, corresponding to antisymmetric spin susceptibilities, we explicitly validate this behavior by examining the electronic band structure and DOS as in the preceding analyses. Figure 5(b) offers further insight by displaying the electronic DOS for $\beta=0$ and $g=0.2$ eV/Å, along with the corresponding electronic band structure $\mathcal{E}_{\vec{k}}$ along the high-symmetry directions $Y\leftarrow\Gamma\to X$ (inset).
As illustrated, the DOS along both the $X$ and $Y$ directions exhibits identical features at $\beta=0$, confirming that in the absence of altermagnetic order, the induced spin Edelstein (de)polarization remains highly isotropic. By contrast, the introduction of a finite altermagnetic order parameter $\beta$, breaks this isotropy by inducing anisotropic spin-splitting in the band structure, thereby rendering the spin Edelstein response directionally dependent. This anisotropy is a crucial attribute for spintronic applications, where control over the spin polarization along specific crystallographic directions is often desired.
As $\beta$ increases, the induced spin polarization decreases and the depolarization point shifts slightly toward larger threshold EPC strengths $g_{\rm c}$, indicating that the altermagnetic strength delays the onset of depolarization. This was already confirmed by Fig. 3(a). Physically, this can be attributed to the enhanced band structure induced by the combined effects of EPC and altermagnetic order, which strongly suppresses the intraband and interband contributions to the susceptibility.
@@ IV.2 Phonon-altermagnetic order coupling effect
As the next investigation, we examine how EPC to the altermagnetic order (due to $C_{4}\mathcal{T}$ symmetry breaking), $\tilde{g}/g$, affects the persistence of spin Edelstein depolarization. It is observed, in Fig. 6(a), that for $\tilde{g}/g<1$, the rate of polarization suppression depends only weakly on $\tilde{g}/g$. Larger values of the ratio (but still below 1) lead to a slower increase in $\chi_{xy}/\chi_{0}$ and slightly shift the zero to larger $g$, implying that a stronger coupling to the staggered potential slightly enhances the susceptibility renormalization.
At $\tilde{g}/g=1$, there is a strong shift of the bands, such that a much larger EPC is required to suppress the spin Edelstein polarization, as the intraband and interband transitions still contribute significantly to the susceptibility. This, in turn, means that stronger $\tilde{g}/g$ stabilizes the transverse polarization more efficiently.
@@
We further highlight that in the absence of explicit EPC to the staggered potential, depolarization still occurs. However, in Fig. 6(b) at $g=0.2$ eV/Å, our DOS analysis and band structure (inset of Fig. 6(b)) reveal that the induced polarization and depolarization remain isotropic at $\tilde{g}/g=0$, since the DOS is identical along the $X$ and $Y$ directions of the Brillouin zone. The parameters $\tilde{g}$ and $\beta$ are then intrinsically linked, such that changes in one necessarily influence the other, reflecting their mutual dependence within the system. Similar to the case of $\beta=0$, such isotropy is undesirable for spintronics, where directional control of spin responses is essential.
IV.3 Rashba spin-orbit coupling effect
@@
IV.4 Electron and hole doping effects
Finally, we examine the case of the doping effect, in which the chemical potential is shifted away from the zero Fermi energy. Figure 8 demonstrates the influence of the chemical potential $\mu$ on the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ as the EPC strength $g$ is varied.
@@ IV.4 Electron and hole doping effects
For positive chemical potentials ($\mu=+0.02$ and $0.04~\mathrm{eV}$), corresponding to electron doping, the behavior shown in Fig. 8(b) differs significantly. The susceptibility curve exhibits a nonmonotonic dependence on $g$, characterized by an intermediate enhancement before eventually decaying. The peak in the susceptibility can be understood in the same way as in Fig. 7(b): for $\mu>0$ and $g>0.15$ eV/Å, an increasing number of states participate in intraband and interband transitions, leading to a temporary increase in susceptibility. Beyond this point, as the number of available states diminishes, the susceptibility begins to decrease. Thus, a strong particle-hole asymmetry appears in the Edelstein response.
@@ V Experimental perspective
Before concluding, we comment on the experimental feasibility of our predictions. A promising route to experimentally probe the proposed effects is to interface a 2D $d$-wave altermagnet—such as thin films of KV2Se2[93], RbV2Te2[94], RuO2 (noting that the altermagnetic character of RuO2 remains under active investigation) [95, 96, 14, 97], and $\kappa$-Cl [98]—with an appropriate substrate at low temperatures to induce lattice vibrations in the altermagnet layer. Crucially, since the underlying phonon-mediated mechanism requires only a linear coupling between lattice displacements and spin states, the proposed effect should be realizable across a broad class of altermagnetic thin films. Electrostatic control over the RSOC can be achieved using top and bottom gate electrodes separated by high-quality dielectric layers. Complementary to phonon generation and gating, angle-resolved photoemission spectroscopy enables direct visualization of spin textures, band splittings, and symmetry-breaking phenomena [99]. Additionally, the Edelstein effect can be probed in Hall-bar geometries through electrical transport measurements.
@@ VI Summary and outlook
In this work, we have examined how slow lattice vibrations shape spin responses in $d$-wave altermagnets, systems that combine vanishing net magnetization with spin-split electronic bands. By formulating the problem within a Rashba continuum model coupled to slow phonons through a Holstein-type interaction, and treating the EPC at the static-Holstein level, we have tracked the evolution of the Edelstein effect over a broad range of parameters. The Kubo formalism provided access to distinguish the specific contributions of intraband and interband processes. In our framework, EPC plays a dual role. It renormalizes the electronic dispersion through an effective shift of the chemical potential and, when tied to a static lattice distortion—such as a piezomagnetically active strain—acts as a symmetry-breaking field that lifts the $C_{4}\mathcal{T}$ symmetry. Thus, EPC is not merely a passive renormalization effect but an active mechanism that reshapes the electronic structure and enables the symmetry-driven phenomena discussed here.
@@
A central result of our study is that increasing EPC gradually suppresses the Edelstein spin polarization, leading to a threshold beyond which the spin polarization collapses to zero, signaling a qualitative restructuring of the spin channels. In this regime, both intraband and interband transitions at the Fermi energy vanish, and the effective Fermi surface, required for the Edelstein response, disappears. The (an)isotropic character of the depolarization strongly depends on altermagnetism: without it, the effect is isotropic, while its interplay with altermagnetism induces pronounced anisotropy and breaks the usual antisymmetry of the spin susceptibilities, highlighting altermagnetism as key for directionally selective spin responses. Our parameter-space analysis shows that the onset of depolarization is tunable: the threshold EPC shifts with staggered lattice coupling, Rashba interaction, and carrier doping. Moving the chemical potential away from zero modifies the threshold, with hole doping accelerating depolarization and electron doping delaying it by enhancing the susceptibility.
@@
As future directions, first, exploring the interplay of phonon-induced depolarization with other spin-orbit interactions, such as Dresselhaus or proximity-induced effects, could reveal richer anisotropic spin responses. Second, extending these studies to multilayer altermagnets or heterostructures may enable enhanced tunability via interlayer coupling and interface engineering. Third, time-resolved experiments, including ultrafast optical or terahertz pump-probe techniques, could probe the dynamical evolution of the Edelstein effect under controlled phonon excitations. Fourth, combining phonon engineering with strain, gating, or substrate manipulation could allow precise control over spin polarization for optimized spintronic devices. Finally, applying these concepts to other materials, such as topological insulators, 2D magnets, or Rashba systems with strong correlations, may reveal universal features of phonon-controlled spin responses. We leave these topics for future work.
@@ Acknowledgments
M. Y. gratefully acknowledges useful discussions with Ulrich Zülicke. M. Y. and J. K. F. were supported by the Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Contract No. DE-FG02-08ER46542 for the formal developments, the numerical work, and the writing of the manuscript. J. K. F. was also supported by the McDevitt bequest at Georgetown University. J. L. was supported by the Research Council of Norway through its Centres of Excellence funding scheme Grant No. 262633 and Grant No. 353894 for writing of the manuscript.
Data availability
The data supporting the findings of this article are openly available in the Zenodo database [100].
Appendix A Derivation of the phonon-dressed low-energy Hamiltonian
In this appendix, we explicitly perform the projection of the lattice electron-phonon Hamiltonian $H_{\mathrm{e\text{-}p}}=\sum_{i,\nu}Q^{\nu}_{i}\,c^{\dagger}_{i}\hat{O}_{\nu}c_{i}$, given in Sec. II.2, onto our low-energy Rashba-altermagnetic subspace in the adiabatic phonon regime, where we restrict to the lowest phonon mode $\nu=0$ at $\vec{q}=\vec{0}$, so that the phonon coordinate is uniform in space, $Q_{0}(\vec{R}_{i})\equiv Q_{0}$. The lattice electron-phonon Hamiltonian then reduces to
$H_{\mathrm{e\text{-}p}}^{(0)}=Q_{0}\sum_{i}\psi_{i}^{\dagger}\left(g_{A}^{x}e_{A}^{(0)x}+g_{B}^{y}e_{B}^{(0)y}\right)\psi_{i}.$ (23)
For the correlated displacement pattern $e_{A}^{(0)x}=1$ and $e_{B}^{(0)y}=1$, the electronic operator associated with the frozen phonon mode is $\hat{O}_{0}=g_{A}^{x}+g_{B}^{y}$. Up to this point, no approximation has been made. Near the $\Gamma$ point, the low-energy electronic Hilbert space is spanned by two bands $s=\pm$ characterized by spin-sublattice (two sublattices $A$ and $B$) locking. Explicitly, the Bloch spinors may be written as $|u_{+}\rangle=(|A,\uparrow\rangle,0)^{T}$ and $|u_{-}\rangle=(0,|B,\downarrow\rangle)^{T}$. These states are orthonormal and satisfy $\sigma_{z}|u_{\pm}\rangle=\pm|u_{\pm}\rangle$.
The projection of $\hat{O}_{0}$ onto the low-energy subspace is
$\hat{O}_{0}=\sum_{s,s^{\prime}=\pm}|u_{s^{\prime}}\rangle\langle u_{s^{\prime}}|\left(g_{A}^{x}+g_{B}^{y}\right)|u_{s}\rangle\langle u_{s}|.$ (24)
Since $g_{A}^{x}$ acts only on sublattice $A$ with spin-up and $g_{B}^{y}$ only on sublattice $B$ with spin-down, the matrix elements are diagonal, $\langle u_{s^{\prime}}|\hat{O}_{0}|u_{s}\rangle=\delta_{ss^{\prime}}g_{\uparrow}^{x}$ for $s=+$ and $\langle u_{s^{\prime}}|\hat{O}_{0}|u_{s}\rangle=\delta_{ss^{\prime}}g_{\downarrow}^{x}$ for $s=$-, where
$g_{\uparrow}^{x}\equiv\langle A,\uparrow|g_{A}^{x}|A,\uparrow\rangle,\qquad g_{\downarrow}^{y}\equiv\langle B,\downarrow|g_{B}^{y}|B,\downarrow\rangle.$ (25)
Thus, in the $\{|u_{+}\rangle,|u_{-}\rangle\}$ basis, we obtain
$\hat{O}_{0}=\begin{pmatrix}g_{\uparrow}^{x}&0\\ 0&g_{\downarrow}^{y}\end{pmatrix}=\frac{g_{\uparrow}^{x}}{2}\left(\sigma_{0}+\sigma_{z}\right)+\frac{g_{\downarrow}^{y}}{2}\left(\sigma_{0}-\sigma_{z}\right),$ (26)
leading to $\hat{O}_{0}=\frac{g_{\uparrow}^{x}+g_{\downarrow}^{y}}{2}\,\sigma_{0}+\frac{g_{\uparrow}^{x}-g_{\downarrow}^{y}}{2}\,\sigma_{z}$. Thus, the resulting low-energy electron-phonon Hamiltonian reads
$H_{\mathrm{e\text{-}p}}=Q_{0}\sum_{i}c_{i}^{\dagger}\left(\frac{g_{\uparrow}^{x}+g_{\downarrow}^{y}}{2}\,\sigma_{0}+\frac{g_{\uparrow}^{x}-g_{\downarrow}^{y}}{2}\,\sigma_{z}\right)c_{i}.$ (27)
In the pristine lattice, the two sublattices $A$ and $B$ are related by a lattice symmetry operation (rotation or mirror, depending on lattice type) followed by time reversal. The degeneracy at $\Gamma$ is therefore protected by the $C_{4}\mathcal{T}$ symmetry. A finite staggered coupling $g_{\uparrow}^{x}\neq g_{\downarrow}^{y}$ requires a lattice distortion that breaks $C_{4}\mathcal{T}$ symmetry. Appropriate symmetry-lowering piezomagnetism strain patterns generate $g_{\uparrow}^{x}\neq g_{\downarrow}^{y}$ [65, 66, 67, 68, 69, 70, 71, 66, 72].